Two conductors of same material of radii r and 2r are connected in 1)series 2)Parallel.Find the ratio of the drift velocities in each case??

Grade:12

2 Answers

Khimraj
5 years ago
​i) as current will be same
 so I = n*e*A1*vd1 =n*e*A2​*vd2
so vd1/vd2 = r12/r22 = 1/4
Hope it clears.
Khimraj
5 years ago
sorry
so vd1/vd2 = r22/r12 = 4/1

?ii) R1/R2 = 4/1 so I1/I2 = 4/1
then vd1/vd2 = (I1/I2)(r22/r12) = 4*4/1*1 = 16/1
Hope it clears.